From 4 red, 5 green, and 6 yellow apples, how many selections of 9 apples are possible if 3 of each color are to be selected?


1
Expert's answer
2021-10-28T07:40:20-0400

"N=C_4^3*C_5^3*C_6^3=4*10*20=800."